Using the Cauchy โ€“Riemann equations verify the following is analytic or not


i) ๐‘ฅ^2 โˆ’ ๐‘ฆ^2 + 2๐‘–๐‘ฅ๐‘ฆ


ii) ๐‘ฅ^2 + ๐‘ฆ^2 โˆ’ 2๐‘–๐‘ฅ๐‘ฆ

1
Expert's answer
2022-03-19T02:40:17-0400

"i:\\u\left( x,y \right) =x^2-y^2\\v\left( x,y \right) =2xy\\\frac{\partial u}{\partial x}=2x,\frac{\partial u}{\partial y}=-2y\\\frac{\partial v}{\partial x}=2y,\frac{\partial v}{\partial y}=2x\\\frac{\partial u}{\partial x}=\frac{\partial v}{\partial y},\frac{\partial u}{\partial y}=-\frac{\partial v}{\partial x}\Rightarrow f\,\,is\,\,analytic\\ii:\\u\left( x,y \right) =x^2+y^2\\v\left( x,y \right) =-2xy\\\frac{\partial u}{\partial x}=2x\\\frac{\partial v}{\partial y}=-2x\\\frac{\partial u}{\partial x}\ne \frac{\partial v}{\partial y}\Rightarrow f\,\,is\,\,not\,\,analytic"